6. What is the minimum work needed to push a 950 kg car 810 m up along a 9.0° incline?
a) Ignore friction
b) Assume the effective coefficient of friction retarding the car is 0.25
1
Expert's answer
2020-05-24T17:59:29-0400

We have given,

mass of the car "m=950Kg" ,inclined distance "d=810m" and the angle of inclination "\theta=9.0^{\circ}" .

a).Let's draw the FBD of the system,


clearly from the above free body diagram (FBD), to lift up the car along the inclined plane,

minimum applied force must be,


"F_{applied}=mg\sin(\theta)"

Thus, work done by the "F_{applied}" will be,


"W_{applied}=F_{applied}\cdot d"

Hence,

"W_{applied}=mgd\sin(\theta)\\
\implies W_{applied}=950\times 9.8\times810\times \sin(9.0^{\circ})\\
\implies W_{applied}=11,79,687.94J"

b).In this case frictional force is acting and coefficient of friction "\mu=0.25" given.

Again draw the FBD of the system,


From the above FBD, we get that the minimum applied force need to be,


"F_{applied}=mg\sin(\theta)+f"

Since,

"f=\mu N"

where, "N" is the normal force acting between incline plane and the given car,thus

"N=mg\cos(\theta)\implies f=\mu mg\cos(\theta)"

Therefore,

"F_{applied}=mg(\sin(\theta)+\mu \cos(\theta))\\
\implies W_{applied}=F_{applied}\cdot d\\
\implies W_{applied}=mgd(\sin(\theta)+\mu \cos(\theta))\\
\implies W_{applied}=950\times9.8\times810\times(\sin(9.0^{\circ})+0.25\cos(9.0^{\circ}))\\
\implies W_{applied}=30,41,752.08J"

Therefore we are done.